# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v1

@@ 6 Conclusion
This paper investigates modality preference in multimodal large language models (MLLMs). We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference. We observe that all 18 tested MLLMs exhibit overall modality bias, and modality preference directions are encoded in MLLMs’ representation space. Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference. Experiments show that the proposed method generalizes well to downstream tasks, such as hallucination mitigation and multimodal machine translation.
@@ A.2 Social Impacts
A.3 Ethical Considerations
All images and generated text contexts, which we use to construct MC2, strictly follow guidelines designed to exclude any harmful, unethical, or offensive content. Human annotators are explicitly instructed to refrain from producing any personally identifiable information or inappropriate language during the annotation process. Furthermore, our benchmark does not involve any comparisons of harmful, ethical, or offensive content between image pairs.
@@ A.4 License of Assets
All images in MC2 are publicly available from COCO [26]. We release our benchmark under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
@@ B.1 Conflict Text Context Generation
To ensure reproducibility and transparency, we include the exact prompts used in our data generation process. These prompts were designed to generate the candidate textual contexts and corresponding answers using GPT-4o-mini [21] and DeepSeekV3 [28]. Below, we provide representative examples of the prompts used during dataset construction given the caption of an image, question, the answer for the question based on image and the task type for the question. For the full list of prompts, please refer to the project repository.
@@
Although the text contexts and answers generated by strong LLMs—filtered through judge MLLMs such as LLaVA1.5-7B [29] and QwenVL-7B [6]—generally yield reliable results, we further incorporate manual inspection to ensure the high quality of data annotations. Specifically, we verify that the visual and textual contexts are indeed in conflict, and that each modality independently supports the corresponding answer to the given question. This involves a two-stage manual review process:
  • Modality-Answer Alignment. First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question. This step is particularly important because the original VQA answers in the TDIUC [22] dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
@@ C.1 Evaluatiion Detail for Modality Preference
We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B [29], LLama3.2-11B-Vision-Instruct [13], LLaVA-OneVision-7B/72B [24], CogVLM2-19B [17], mPLUG-Owl3-24-07 [49], Qwen2VL-7B [43], GLM-4V-9B [10], SPHINX-V2-1K [27], InternVL3-9B/14B/38B/78B [2], LLaVA-next-7B/13B/34B [29] and Qwen2.5VL-7B/32B/72B [7]. All the open-source models are evaluated using NVIDIA A100 or A800 GPUs. We also evaluate the proprietary model, GPT-4o-mini [21] via the official API.
@@
We provide the results of modality preference for several models in Table 1 in the main text. More detailed modality preference evaluation results are presented in Table 7. We evaluated 18 models, including 17 open-source models and the commercial model, GPT4o-mini.
C.2 The Details for Controlling Modality Preference
Noisy image and text grammar error generation
For noisy images, we initially adopte two approaches: adding Gaussian noise to the image and applying a uniform linear motion blur kernel. In our experiments, both methods yield similar results in terms of model performance, indicating comparable effects on the evaluation. We report the results based on the Gaussian noise method in this paper.
To generate grammatically incorrect text, we employ DeepSeek-V3 to rewrite the original text context. We find that DeepSeekV3 performs well in introducing localized grammatical errors while preserving the overall structure. We use two different prompting strategies, both of which produced similar results. In this paper, we report the results based on the V1 version of grammar error generation. The specific prompt used is shown below:
Text Context Grammar Error Generation (V1) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes. These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words. Despite these errors, the core meaning of the sentence must remain unchanged. The text to be rewritten is: {original text context}. Output the revised text enclosed in <context2> </context2> tag. Only output the revised text and don’t explain anything.
Text Context Grammar Error Generation (V2) using DeepSeek-V3 Instruction: # Please rewrite the following text, intentionally introducing three grammar errors and spelling mistakes. These errors should include, but are not limited to, subject-verb agreement mistakes, incorrect use of tenses, improper punctuation, misplacement of modifiers, and misspelled words. Despite these errors, the core meaning of the sentence must remain unchanged. Additionally, do not alter the meaning related to the provided critical word. The text to be rewritten is: {original text context}. the critical word is: {answer based on text context}. Output the revised text enclosed in <context1> </context1> tag. Only output the revised text and don’t explain anything.
The detail for controlling modality preference using noise data
We first test the model’s ability to understand a single noisy modality (either a noisy image context or a text context with grammar errors). As shown in Table 13 and Table 12, the model’s understanding of the single-modality content remains close to 95% for the noisy image context and close to 98% for the noisy text context, which is nearly the same as its performance on the original single-modality content (shown in Table 11 and Table 10). This rules out the possibility that the shift in modality preference due to the added noise is caused by a decrease in the model’s understanding of the single modality. Instead, it suggests that MLLMs may implicitly associate modality preference with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning.
We provide the results of modality preference using either the noisy image context or the text context with grammer error in Figure 4 in main text. We also provide more results for more models in Table 14 and Table 15.
@@ C.2 The Details for Controlling Modality Preference
In Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B and Qwen2.5VL-7B. We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 8 and Table 9. For each setting, we report the probability of responses exhibiting a text preference (measured by vision-based accuracy) and the probability of responses exhibiting a vision preference (measured by text-based accuracy).
C.3 The details of PCA Analysis
In Section 4.2, we mentioned the PCA analysis details regarding the "Modality Preference Direction in Representation Space." Here, we provide a more in-depth description of the setup. We extract the model’s hidden representations from the last token of the input across different layers. Then, we apply the PCA method to reduce the dimensionality to two dimensions for visualization. The following settings were visualized:
@@ D.1 Details for Pattern of hidden states
In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B. As shown in Figure 11, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-14B.
@@
For each model, we selected layers with large absolute differences and small standard deviations. This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
@@
Figure 11: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B models from left to right.
D.2 More Results for Steering Models Toward Vision or Text Modality Preference
In the main text, we present the results of steering Qwen2VL-7B and Qwen2.5VL-7B towards vision and text modality preferences, respectively. Here, we provide additional results on modality preference control, including experiments with InternVL3-14B and LLaVA-OneVision-7B. As shown in Table 6, the proposed method significantly outperforms the baseline in both directions of modality preference.
D.3 Downstream Task Applications
By steering the model’s modality preference towards the vision modality, we reduce vision hallucinations in MLLMs. Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation. This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output.
Details of Vision Hallucination Mitigation
PhD [30] is a multimodal hallucination evaluation benchmark that increases the risk of hallucination by introducing "hitem" items, intentionally guiding the model to produce hallucinated words, thereby increasing the difficulty of the model’s responses. Phd-icc and Phd-ica are subsets of PhD, which respectively increase hallucination risks by introducing incorrect text context and inaccurate text context. For testing convenience, we randomly selected 1,000 samples from the original Phd-cc and Phd-ica datasets for evaluation.
By steering modality preference towards the vision context, the model is encouraged to focus more on visual information, which helps reduce vision hallucinations. As shown in Table 3, the proposed method outperforms most baseline methods across nearly all subcategories of the Phd-icc and Phd-ica datasets. Furthermore, a detailed analysis of the two datasets reveals that most samples in Phd-iac introduce only irrelevant text context, whereas Phd-icc introduces incorrect text context, which makes it easier for the model to generate hallucinations. Proposed method significantly outperforms the baseline methods on Phd-icc, indicating that the proposed method is more effective in mitigating hallucinations.
Details of Multimodal Multimodal machine translation
@@ D.3 Downstream Task Applications
Ambigcaps [25] benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets. We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality. We also use our method to steer the modality preference towards the text modality. As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish.
Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
Vision MLLM-only 32.0 21.6 2.8 24.8 2.4 30.0 11.6 42.4 20.9
Inst Design 55.6 31.2 12.0 30.8 3.2 36.4 22.4 61.2 31.6
ours 75.6 40.8 18.0 38.0 10.0 53.2 32.0 74.4 42.8
Text MLLM-only 8.8 48.4 60.0 52.0 47.2 21.2 52.8 9.6 37.5
Inst Design 44.8 68.4 82.4 54.8 43.2 50.0 58.8 19.6 52.8
ours 49.6 71.2 85.2 58.0 47.6 48.4 68.4 58.4 57.1
Table 6: Performance by steering LLaVA-OneVision-7B towards vision and InternVL3-7B towards text.
@@
OneVision-7B 98.0 95.2 100.0 97.2 97.2 97.2 98.0 99.2 97.8
Qwen2.5VL-7B 98.8 96.4 100.0 99.2 97.6 97.6 98.8 99.2 98.4
CogVLM2-19B 98.0 94.4 99.2 96.0 91.2 97.6 97.6 99.6 96.7
InternLM-XC2.5-7B 97.6 97.2 99.2 99.2 97.6 97.6 98.8 98.8 98.2
GLM-4V-9B 98.8 96.4 99.6 97.2 98.8 98.4 99.6 100.0 98.6
SPHINX-V2-1K 98.4 97.6 99.2 98.8 97.6 98.4 98.4 99.6 98.5
LLaVA1.6-7B 98.0 90.8 98.8 96.8 96.8 97.2 96.0 96.4 96.4
Table 12: Accuracy of question answering in the MC2 dataset when only unimodal noisy textual context is provided.
OneVision-7B 98.4 95.6 94.0 98.4 78.4 99.2 95.6 98.0 94.7
Qwen2.5VL-7B 98.8 96.4 100.0 99.2 99.2 100.0 100.0 98.8 99.3
CogVLM2-19B 99.6 97.6 87.6 94.4 91.6 98.8 97.2 98.48 95.7
InternLM-XC2.5-7B 98.4 96.0 94.4 94.0 95.2 100.0 97.6 95.2 96.4
GLM-4V-9B 99.2 98.4 95.6 98.4 98.4 99.6 100.0 99.2 98.6
SPHINX-V2-1K 98.4 97.2 90.0 98.8 98.4 97.6 96.8 98.0 96.9
LLaVA1.6-7B 98.4 94.8 92.4 93.6 84.0 96.8 93.6 95.6 93.7
Table 13: Accuracy of question answering in the MC2 dataset when only unimodal noisy visual context is provided.
OneVision-7B 42/29.2 31.2/41.6 4.4/88 31.6/44 4.4/82 39.2/29.6 16.4/66.8 50.8/26 27.5/50.9
Qwen2.5VL-7B 74.0/12.0 53.20/36.40 24.4/62.4 58/29.2 54.8/35.2 78/16 52.8/35.6 82.4/12.4 59.7/29.9
InternLM-XC2.5-7B 75.6/18.4 32.8/57.2 56.0/36.0 26.4/59.2 11.6/78.0 61.2/28.0 52.8/36.4 74.0/20.0 48.8/41.7
GLM-4V-9B 48.0/31.6 37.6/48.8 10.0/81.2 34.0/56.4 18.8/70.8 64.8/27.2 33.6/55.6 61.6/25.2 38.6/49.6
SPHINX-V2-1K 46.0/44.8 18.0/77.6 1.2/98.0 21.6/70.8 11.6/80.8 30.0/64.8 30.8/60.0 62.8/28.8 27.8/65.7
CogVLM2-19B 46.0/39.6 34.0/47.2 12.4/68.4 22.8/52.0 8.4/70.0 37.2/39.2 30.0/55.2 62.8/24.4 31.7/49.5
LLaVA1.6-7B 11.2/71.2 6.4/83.2 1.2/91.6 5.2/76.4 0.4/88.8 7.2/71.6 5.6/70.4 28.8/44.4 8.3/74.7
Table 14: Accuracy of question answering in the MC2 dataset when both noisy textual context and visual context are provided but the instruction does not specify which modality context should be used. Values are reported as vision-based accuracy/text-based accuracy for each model.
OneVision-7B 24/44 16.4/66.4 2/96.8 17.2/66.4 0.80/87.2 20.4/47.2 8/78.8 35.6/42.8 15.5/66.2
Qwen2.5VL-7B 48.8/23.6 25.6/64.8 6/88 33.2/59.6 39.2/54.8 58.8/32.8 30/56.8 68/56.8 38.7/50.3
InternLM-XC2.5-7B 52.8/33.6 15.2/76.8 24.4/64.4 12.8/78.4 3.6/90.4 33.2/56.8 24.4/64.8 54.8/37.2 27.7/62.8
GLM-4V-9B 32.4/50.4 19.6/73.2 4.0/91.6 21.2/69.2 11.2/78.4 41.6/46.8 15.2/76.8 41.2/42.0 23.3/66.1
SPHINX-V2-1K 29.2/58.0 11.2/86.4 0.8/98.8 16.8/80.4 5.6/90.4 18.0/76.4 18.0/76.0 52.8/40.4 19.1/75.9
CogVLM2-19B 32.0/45.2 22.0/60.0 3.6/84.8 14.0/60.8 6.4/73.2 18.8/49.2 18.8/65.6 46.0/36.4 20.2/59.4
LLaVA1.6-7B 8.4/79.2 4.0/87.2 0.4/94.8 3.2/82.4 0.4/92.0 4.4/77.6 4.4/76.8 22.8/50.8 6.0/80.1
Table 15: Accuracy of question answering in the MC2 dataset when both textual and noisy visual contexts are provided but the instruction does not specify which modality context should be used. Values are reported as vision-based accuracy/text-based accuracy for each model.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. 2.
    Limitations
  5. Question: Does the paper discuss the limitations of the work performed by the authors?
  6. Answer: [Yes] See discussions in Suppl.
  7. 3.
    Theory assumptions and proofs
  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  9. Answer: [N/A]
  10. 4.
    Experimental result reproducibility
  11. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  12. Answer: [Yes] See details in Suppl.
  13. 5.
    Open access to data and code
  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  15. Answer: [Yes] See details in Suppl.
  16. 6.
    Experimental setting/details
  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  18. Answer: [Yes] See details in Suppl.
  19. 7.
    Experiment statistical significance
  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  21. Answer: [Yes] See details in Suppl.
  22. 8.
    Experiments compute resources
  23. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  24. Answer: [Yes] See details in Suppl.
  25. 9.
    Code of ethics
  26. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  27. Answer: [Yes]
  28. 10.
    Broader impacts
  29. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  30. Answer: [Yes]
  31. 11.
    Safeguards
  32. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  33. Answer: [Yes] See details in Suppl.
  34. 12.
    Licenses for existing assets
  35. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  36. Answer: [Yes]
  37. 13.
    New assets
  38. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  39. Answer: [Yes] See detail in Suppl.
  40. 14.
    Crowdsourcing and research with human subjects
  41. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  42. Answer: [Yes] See details in Suppl.
  43. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  44. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  45. Answer: [Yes] See details in Suppl.
  46. 16.
    Declaration of LLM usage
  47. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  48. Answer: [Yes] See details in Suppl.

### v2

@@ 6 Conclusion
This paper investigates modality preference in multimodal large language models (MLLMs). We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference. Besides, we find the direction of modality preference can be captured within the latent representations of MLLMs. Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference. Experiments show that the proposed method generalizes well to downstream tasks, such as multimodal machine translation and multimodal understanding tasks.
@@ A.2 Social Impacts
A.3 Use of LLM
In this work, we use the LLMs including GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a), and MLLMs including LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023) to help annotate the text context in MC2, as detailed in Section B. We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024). Besides, we also utilize the LLMs to correct the grammatical errors.
@@ A.4 License of Assets
All images in MC2 are publicly available from COCO (Lin et al., 2014). We release our benchmark under a Creative Commons Attribution 4.0 License (CC BY 4.0) to enhance global accessibility and foster innovation and collaboration in research.
@@ B.1 Conflict Text Context Generation
To ensure reproducibility and transparency, we include the exact prompts used in our data generation process. These prompts were designed to generate the candidate textual contexts and corresponding answers using GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a). Below, we provide representative examples of the prompts used during dataset construction given the caption of an image, question, the answer for the question based on image and the task type for the question. For the full list of prompts, please refer to the project repository.
@@
Although the text contexts and answers generated by strong LLMs—filtered through judge MLLMs such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023)—generally yield reliable results, we further incorporate manual inspection to ensure the high quality of data annotations. Specifically, we verify that the visual and textual contexts are indeed in conflict, and that each modality independently supports the corresponding answer to the given question. This involves a two-stage manual review process:
  • Modality-Answer Alignment. First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question. This step is particularly important because the original VQA answers in the TDIUC (Kafle & Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
@@ C.1 Evaluatiion Detail for Modality Preference
We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), LLaVA-OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025). All the open-source models are evaluated using NVIDIA A100 or A800 GPUs. We also evaluate the proprietary model, GPT-4o-mini (Hurst et al., 2024a) via the official API.
@@
We provide the results of modality preference for several models in the left panel of Figure 2 in the main text. More detailed modality preference evaluation results are presented in Table 6.
C.2 Can the Vision Ratio provide guidance for downstream task performance?
We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multimodal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022). We compute the average score on all datasets, where MME score is normalized between 0-1, as shown in Table 5.
Model MMMU MME MMBench RealworldQA MMStar HallBench InfoVQA ChartQA Avg Vision Ratio
Qwen2.5VL-7B 58.6 83.8 83.5 68.5 63.9 52.9 82.6 87.3 75.5 59.6
Qwen2.5VL-72B 70.2 87.4 88.6 75.7 70.8 55.2 87.3 89.5 81.4 78.6
InternVL3-9B 57.7 84.7 83.4 70.5 66.3 51.2 79.6 86.2 75.5 41.2
InternVL3-14B 67.1 88.5 85.6 70.7 68.8 55.1 83.6 87.3 78.8 55.0
InternVL3-38B 70.1 90.1 87.6 75.6 71.5 57.1 85.0 89.2 81.3 62.4
InternVL3-78B 72.2 91.1 89.0 78.0 72.5 59.1 86.5 89.7 82.7 81.5
LLaVA-OneVision-7B 47.9 71.2 83.2 66.3 61.7 31.6 68.8 80.0 68.4 26.3
LLaVA-OneVision-72B 55.7 80.8 85.8 71.9 65.8 49.0 74.9 83.7 74.1 30.1
Qwen2VL-7B 54.1 83.1 83.0 70.1 60.7 50.6 76.5 83.0 72.9 16.3
LLaVA-Next-7B 37.6 63.2 69.2 57.8 37.6 27.6 31.6 51.9 49.8 8.5
LLaVA-Next-13B 37.3 62.3 70.0 57.6 40.4 31.8 34.9 59.0 51.6 9.7
LLaVA-1.5-7B 35.7 64.6 69.2 54.8 33.1 27.6 22.4 17.8 42.5 13.4
LLaVA-1.5-13B 37.0 63.6 66.5 55.3 34.3 24.5 24.9 18.5 42.9 15.0
Table 5: Performance comparison across benchmarks for different models measured by accuracy (%) and Vision Ratiio score (%).
C.3 The Details for Controlling Modality Preference
@@
In the left panel of Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B. We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 7 and Table 8. For each setting, we report the results measured by $S_{vision}$, vision-based accuracy and $S_{text}$, text-based accuracy.
The details of PCA Analysis
In Section 4.4, we use the PCA analysis regarding the Modality Preference Direction in Representation Space. Here, we provide a more detailed description of the setup. We extract the model’s hidden representations from the last token of the input across different layers. Then, we apply the PCA method to reduce the dimensionality to two dimensions for visualization. The following settings were visualized:
@@ D.1 Details for Pattern of hidden states
In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B. As shown in Figure 10, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-8B. For each model, we selected layers with large absolute differences and small standard deviations. This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
@@
Figure 10: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
D.2 Evaluation of visual understanding and multimodal multimodal machine translation
PhD (Liu et al., 2024c) is a visual understanding benchmark and includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination. For testing convenience, we randomly selected 1,000 samples from the original Phd-cc and Phd-ica datasets for evaluation. By steering the model’s modality preference toward the vision modality, we strengthen its visual understanding ability and mitigate vision hallucinations in MLLMs.
@@ D.2 Evaluation of visual understanding and multimodal multimodal machine translation
Ambigcaps (Li et al., 2021) benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets. We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality. Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multimodal machine translation. This adjustment prevents the model from over-relying on visual content and from introducing spurious objects or extraneous details into the translation output.